OpenStudy (anonymous):

how do i simplify 2x^2/7x2*21y^11/5x^9

OpenStudy (id21):

You must use brackets for better understanding

OpenStudy (anonymous):

(2x^2/7x^2)*(21y^11/5x^9)

OpenStudy (anonymous):

do I just make the reciprocal in the second pair of brackets

OpenStudy (angela210793):

w8 sorry..I've made smth wrong

OpenStudy (anonymous):

ok well any help on this problem will be great

OpenStudy (angela210793):

\[2x^2/7x^2*21y^11/5x^9=4y^11/35x^9\]

OpenStudy (anonymous):

thanks that really helped

OpenStudy (anonymous):

how were you wrong

OpenStudy (id21):

\[\frac{2x^2}{7x^2}\times \frac{21y^{11}}{5x^9}=\frac{2x^2 \times x^{-2}\times x^{-9}\times 21y^{11}}{7\times 5}=\frac{2x^{2-2-9}\times21y^{11}}{7\times5}=\frac{6}{5}x^{-9}y^{11}=\frac{6}{5}\frac{y^{11}}{x^9}\]